OpenStudy (anonymous):

Is this correct ? Factor: 9x2 - 64 (9x + 8)(x - 8)

OpenStudy (callisto):

nope...difference of 2 squares is a^2 - b^2 = (a+b)(a-b) in your case, a=3x, b=8

OpenStudy (anonymous):

so (3x - 8)(3x - 8)

OpenStudy (callisto):

nope... one is +, one is -

OpenStudy (anonymous):

3x + 8)(3x - 8)

OpenStudy (callisto):

yes

Directrix (directrix):

Think of yours as (3x) ^2 - 8^2 --> difference of two squares. Follow the pattern: (3x + 8) (3x - 8)